# Scalable bayesian shadow tomography for quantum property estimation with set transformers

### v1

@@ I Introduction
Quantum technologies promise advances in communication, computation, and sensing by harnessing phenomena such as superposition and entanglement [29, 12, 10]. Because quantum protocols rely heavily on precise quantum state preparation, the ability to accurately characterize quantum states formed in the laboratory or on noisy quantum processors is central to the progress of these technologies. The exponential growth of a quantum system’s Hilbert space with qubit number renders many existing state tomography protocols infeasible beyond a handful of qubits: reconstructing a full density matrix inevitably requires an exponential amount of computation in the worst case [18, 20, 31].
@@
To address this scalability issue, shadow tomography circumvents full-state reconstruction by focusing on estimating only those scalar properties of a state (mostly linear functions) that are of experimental interest [1, 22]. By sampling randomized measurements and computing simple linear inversion estimators, classical shadows provide accurate estimates of expectation values and certain non-linear functions using few measurements. Nonetheless, these estimators remain non-Bayesian, often exhibiting high variance for a finite number of measurements and lacking a principled way to incorporate prior knowledge about the state.
@@
This work investigates Bayesian shadow tomography, combining motivations from Bayesian quantum state estimation with the classical shadows protocol. In shadow tomography, the goal is not to reconstruct the entire quantum state. Instead, it focuses on estimating specific properties or “shadows” of the state directly. Our objective is to extend this paradigm to the Bayesian estimation setting. Although Bayesian approaches to quantum tomography are present in the literature, the body of work remains relatively limited [7, 34, 15, 27]. Most existing methods focus on reconstructing the quantum state itself, which inherently requires an exponential amount of data and computation. To the best of our knowledge, we are the first to propose a Bayesian and scalable machine learning approach that directly estimates scalar functions of an unknown quantum state without performing full state reconstruction. This is achieved by efficiently encoding each of the $N$ measurement outcomes into a fixed-dimensional feature vector, so that the model input grows linearly in $N$ and only polynomially in the qubit number nn—namely $O(n)$ for random Pauli measurements and $O(n^{2})$ for random Clifford measurements. This encoding makes it possible to train permutation-invariant (PI) networks that scale to large $N$ and large $n$ without explicit reference to the $2^{n}$-dimensional Hilbert space (see FIG. 1). The use of PI models to account for the unordered nature of random measurement data has proven effective in classification tasks [37]. However, our experiments suggest that their direct extension to regression tasks yields limited performance. We suspect this difficulty arises because shadow inversion involves evaluating parity-like functions over measurement outcomes—an extremely rugged mapping for neural networks to capture. Instead, we formulate our estimator via residual learning, so that the network only needs to learn a correction term atop a fixed baseline.
@@
Our method employs a non-adaptive measurement strategy, meaning the measurement settings are fixed in advance and not adjusted based on previous measurement outcomes. We focus on the post-processing or final estimation stage of an experiment that has already been conducted, rather than on adaptively collecting data. The classical shadows protocol itself is essentially a form of non-adaptive linear inversion tomography [35, 19, 22]. Moreover, the shadow estimation procedure operates without incorporating any prior information about the state, and individual snapshots of the state can even correspond to non-physical density operators. In other words, classical shadow tomography does not perform a Bayesian update. Nevertheless, any data obtained from measurements can be used to update a prior distribution over the state. Therefore, inspired by the success of classical shadow tomography, we propose to leverage measurement statistics and shadow estimation algorithms within a Bayesian tomography framework.
@@ II.2 Prior works
Quantum state tomography (QST) aims to reconstruct the quantum state or infer its properties from measurement data. Traditional Bayesian approaches to QST update a prior distribution over the density operator using observed outcomes. Such methods yield principled uncertainty quantification but often suffer from computational cost [4, 15, 27]. A recent effort has mitigated this challenge via neural-network–enhanced Bayesian inference [26], but it remains focused on full state reconstruction.
Parallel to Bayesian methods, non‑Bayesian machine‑learning techniques have been developed to estimate scalar functions of quantum states directly. LLM4QPE [36] is an autoregressive model that treats quantum property estimation as a language task to predict various properties. QUADIM [38] adopts a diffusion-based approach for the same purpose. However, these models assume access to a classical description of the system’s evolution and do not enforce permutation invariance of measurement records.
Another line of work employs neural networks to accelerate full-state QST through adaptive measurement strategies or compressed representations [40, 33, 24]. More recently, transformer-based architectures have been used in reconstructing density matrices from measurement data [42, 32, 41, 28]. While these methods achieve remarkable accuracy, they often focus on recovering the density matrix rather than directly targeting scalar quantities. This incurs unnecessary computational overhead when only a few properties are of interest.
This work develops a Bayesian inference framework enhanced by residual learning and implemented with a PI neural network architecture, aiming to estimate scalar functions of an unknown quantum state from classical measurement data. Rather than reconstructing the full density matrix, our method focuses on efficient, scalable estimation of quantities like fidelity and Rényi entanglement entropy. It exploits the i.i.d. nature of non-adaptive measurements to design an architecture that generalizes across varying system sizes and numbers of measurements.
@@ III Problem setup
In a fully Bayesian approach, one would instead seek the posterior expectation of $f(\rho)$ given the observed data. That is, the ideal Bayesian estimator is the conditional expectation of $f(\rho)$ under the posterior distribution of $\rho$ after observing all $N$ measurement outcomes. Computing this exact posterior expectation is generally intractable for large quantum systems due to two fundamental challenges: first, it requires performing continuous integrations over state spaces; second, the dimension of the quantum state $\rho$ grows exponentially with the number of qubits. In the next section, we describe our learning-based framework for approximating the Bayesian estimator, which enables the incorporation of arbitrary priors without explicitly evaluating high-dimensional integrals.
@@ IV.1 PI network and residual learning
the Bayesian posterior expectation of $f(\rho)$ using the measurement outcomes as input. Instead of directly predicting $f(\rho)$, we find it advantageous to leverage any existing estimator $g$ (such as the classical shadows estimator) and train the network to predict the residual error relative to this estimator. In other words, the network learns to predict the difference between the current estimator’s output $\hat{f}(\rho)$ and the true value $f(\rho)$. By correcting this difference, the network effectively learns to output a refined estimate closer to the Bayesian ideal. That is, we train the model to predict the expected estimation error conditioned on all $N$ measurement results.
Even if each measurement choice can depend on all previous outcomes, the final posterior depends only on the multiset of observed values. Formally,
@@
Our model is specifically designed to capture this permutation invariance. We employ the set transformer architecture [25], a neural network tailored for set-input problems, to process the collection of measurement outcome features. The set transformer uses a self-attention mechanism to produce a PI representation of the input set. They have been shown to be universal approximators for PI functions on sets. Following its original implementation, the network processes the input
@@
$\sigma(x,F)=F_{\text{c}}(1-\tanh|x|)+\mathbf{1}_{\mathbb{R}^{+}}(x)\tanh x,$
where $F_{\text{c}}=\min(\max(F,f_{\text{u}}),f_{\text{l}})$ is a clipped estimate and $\mathbf{1}_{A}(x)=[x\in A]$ is the indicator function. At $x=0$, we have $\sigma(0,F)=F_{\text{c}}$.
Alternatively, one could learn the estimator directly in the form $\text{Sigmoid}(x)$, rather than using a residual correction. However, as shown in the experimental results, this direct approach is outperformed by classical shadows as the number of measurements $N$ increases. This is because inverting the shadow requires learning parity-like functions that conventional learning algorithms notoriously struggle to approximate [3, 14, 39, 11].
@@ IV.3 Encoding of measurement statistics
Each Clifford circuit is specified by its stabilizer tableau of size $2n\times(2n+1)$ [13, 2, 5]. Concretely, each of the $2n$ rows consists of:
@@ IV.4 Application to fidelity estimation
As a first demonstration, we consider direct fidelity estimation (DFE), the task of estimating the fidelity of the unknown state $\rho$ with a target entangled state. Specifically, we take the target state to be the $n$-qubit Greenberger-Horne-Zeilinger (GHZ) state [16, 17] defined as
@@
The fidelity with the GHZ state is $f(\rho)=\langle\psi_{\mathrm{GHZ}}|\rho|\psi_{\mathrm{GHZ}}\rangle$. There exists an efficient procedure for estimating the fidelity of an unknown quantum state with the GHZ state [8]. Algorithm 2 in Appendix G describes this procedure. This baseline procedure is referred to as the “Shadow” estimator for fidelity. The estimator in Algorithm 2 has the following desirable properties.
@@ IV.5 Application to entanglement entropy estimation
As a second example, we consider the task of estimating a non-linear function of the state: the second-order Rényi entanglement entropy across a bipartition of the qubits [21, 23, 6]. Let the qubits be partitioned into subset $A$ and its complement $B$ (for instance, $A$ could be the first $n/2$ qubits and $B$ the remaining $n/2$ qubits). The quantity of interest can be expressed as
@@ V.3 Second-order Rényi entanglement entropy estimation for the GHZ state
For Rényi entropy, each instance consists of a random $n$-qubit state (with a chosen bipartition $A|B$) along with $N$ simulated measurement outcomes according to Algorithm 3, and the corresponding ground truth entanglement entropy $f(\rho)=\text{tr}((\rho\otimes\rho)S_{A})$. The input state was the GHZ state, but an adversary randomly replaces it with a spurious state $\omega$, sampled according to the Hilbert-Schmidt metric [43]. The replacement occurred with probability $p$ drawn uniformly from the interval $(0,1)$:
@@ V.4 Performance evaluation
FIG. 6(a) shows that direct learning of the target functional outperforms residual learning for small $N$. In this regime, the variance of the classical shadow estimator is so large that a crude estimate drawn from the feasible range of $f(\rho)$ can beat it. However, as $N$ increases and the shadow estimator becomes sufficiently accurate, residual learning outperforms direct learning. This indicates that approximating the shadow inversion from measurement data to a point estimate of $f(\rho)$ is much more challenging than learning only the bias of an existing estimator, since many shadow inversion steps include evaluations of parity-like functions (see Algorithm 2 and Eq. (11)).
@@ VI Discussion
Our approach inherits a key limitation of data-driven methods: since the network is trained to learn the posterior correction of the classical shadow estimator, it requires a large training set in order to generalize reliably. Each training sample must be labeled with an almost exact value of the target property—i.e. the ground-truth posterior mean—which in practice can only be obtained by performing a very large number of measurements. Thus, although our architecture scales efficiently with the number of qubits and measurements, the cost of generating labels poses an obstacle when the variance of the shadow estimator grows with system size.
The problem of estimating properties of a quantum state in the presence of imperfect control and readout [9] can be understood naturally within a Bayesian inference framework: one treats the parameters characterizing noisy unitaries and measurement channels as learnable latent variables, and then incorporates that information into posterior updates over the true quantum state.
Looking forward, one promising direction is to bypass the intermediate classical shadow reconstruction and instead update a posterior distribution directly over the target functional $f(\rho)$. By framing the problem as inference on the scalar quantity $f(\rho)$, one could potentially further reduce sample complexity and computational overhead. Another avenue for improvement lies in adaptive measurement strategies, where each subsequent measurement choice depends on previous outcomes to maximize information gain about $\rho$ or $f(\rho)$ [33, 24]. Adaptivity can offer performance gains, and permutationally invariant architectures can still be employed to process measurement records.
@@ Appendix B Classical shadows
In the classical shadows protocol [22], one performs randomized measurements on many independent copies of an unknown state $\rho$. Each measurement outcome is used to construct a snapshot $\hat{\rho}$—an unbiased estimator of the state—such that by averaging these snapshots one can estimate any linear function of $\rho$. For example, for a Hermitian observable $O$, the classical shadows estimator satisfies $\mathbb{E}[\text{tr}(\hat{\rho}O)]=\text{tr}(\rho O)$, assuming $\mathbb{E}[\hat{\rho}]=\rho$ by construction.
@@ 2 $\texttt{sum}\leftarrow 0$
Algorithm 1 Classical shadow estimation of $\text{tr}(\rho O)$ [22]
@@
where each $U_{i}$ is independently sampled uniformly at random from the set $\{\mathbb{I},H,HS^{\dagger}\}$. This corresponds to performing local measurements in a randomly chosen basis from $\{X,Y,Z\}$. As will be discussed in the main text, this work considers DFE with Pauli measurements, where the target state is the GHZ state [16, 17]. In such cases, more efficient alternatives to the direct use of (5) are available, and we adopt one of these optimized methods.
@@
where $U$ is sampled uniformly from the $n$-qubit Clifford group $\mathrm{Cl}(2^{n})$ [30].

### v2

@@ I Introduction
Quantum technologies promise advances in communication, computation, and sensing by harnessing phenomena such as superposition and entanglement [montanaro2016quantum, degen2017quantum, cozzolino2019high]. Because quantum protocols rely heavily on precise quantum state preparation, the ability to accurately characterize quantum states formed in the laboratory or on noisy quantum processors is central to the progress of these technologies. The exponential growth of a quantum system’s Hilbert space with qubit number renders many existing state tomography protocols infeasible beyond a handful of qubits: reconstructing a full density matrix inevitably requires an exponential amount of computation in the worst case [gross2010quantum, haah2016sample, o2016efficient].
@@
To address this scalability issue, shadow tomography circumvents full-state reconstruction by focusing on estimating only those scalar properties of a state (mostly linear functions) that are of experimental interest [aaronson2018shadow, huang2020predicting]. By sampling randomized measurements and computing simple linear inversion estimators, classical shadows provide accurate estimates of expectation values and certain non-linear functions using few measurements. Nonetheless, these estimators remain non-Bayesian, often exhibiting high variance for a finite number of measurements and lacking a principled way to incorporate prior knowledge about the state.
@@
This work investigates Bayesian shadow tomography, combining motivations from Bayesian quantum state estimation with the classical shadows protocol. In shadow tomography, the goal is not to reconstruct the entire quantum state. Instead, it focuses on estimating specific properties or “shadows” of the state directly. Our objective is to extend this paradigm to the Bayesian estimation setting. Although Bayesian approaches to quantum tomography are present in the literature, the body of work remains relatively limited [buvzek1998reconstruction, schack2001quantum, granade2016practical, lukens2020practical]. Most existing methods focus on reconstructing the quantum state itself, which inherently requires an exponential amount of data and computation. To the best of our knowledge, we are the first to propose a Bayesian and scalable machine learning approach that directly estimates scalar functions of an unknown quantum state without performing full state reconstruction. This is achieved by efficiently encoding each of the $N$ measurement outcomes into a fixed-dimensional feature vector, so that the model input grows linearly in $N$ and only polynomially in the qubit number nn—namely $O(n)$ for random Pauli measurements and $O(n^{2})$ for random Clifford measurements. This encoding makes it possible to train permutation-invariant (PI) networks that scale to large $N$ and large $n$ without explicit reference to the $2^{n}$-dimensional Hilbert space (see FIG. 1). The use of PI models to account for the unordered nature of random measurement data has proven effective in classification tasks [tang2024ssl4q]. However, our experiments suggest that their direct extension to regression tasks yields limited performance. We suspect this difficulty arises because shadow inversion involves evaluating parity-like functions over measurement outcomes—an extremely rugged mapping for neural networks to capture. Instead, we formulate our estimator via residual learning, so that the network only needs to learn a correction term on the top of a fixed baseline.
@@
Our method employs a non-adaptive measurement strategy, meaning the measurement settings are fixed in advance and not adjusted based on previous measurement outcomes. We focus on the post-processing or final estimation stage of an experiment that has already been conducted, rather than on adaptively collecting data. The classical shadows protocol itself is essentially a form of non-adaptive linear inversion tomography [sugiyama2013precision, guctua2020fast, huang2020predicting]. Moreover, the shadow estimation procedure operates without incorporating any prior information about the state, and individual snapshots of the state can even correspond to non-physical density operators. In other words, classical shadow tomography does not perform a Bayesian update. Nevertheless, any data obtained from measurements can be used to update a prior distribution over the state. Therefore, inspired by the success of classical shadow tomography, we propose to leverage measurement statistics and shadow estimation algorithms within a Bayesian tomography framework.
@@ II.2 Prior works
Quantum state tomography (QST) aims to reconstruct the quantum state or infer its properties from measurement data. Traditional Bayesian approaches to QST update a prior distribution over the density operator using observed outcomes. Such methods yield principled uncertainty quantification but often suffer from computational cost [blume2010optimal, granade2016practical, lukens2020practical]. A recent effort has mitigated this challenge via neural-network–enhanced Bayesian inference [lohani2023demonstration], but it remains focused on full state reconstruction.
Parallel to Bayesian methods, non‑Bayesian machine‑learning techniques have been developed to estimate scalar functions of quantum states directly. LLM4QPE [tang2024towards] is an autoregressive model that treats quantum property estimation as a language task to predict various properties. QUADIM [tangquadim] adopts a diffusion-based approach for the same purpose. However, these models assume access to a classical description of the system’s evolution and do not enforce permutation invariance of measurement records.
Another line of work employs neural networks to accelerate full-state QST through adaptive measurement strategies or compressed representations [torlai2018neural, quek2021adaptive, lange2023adaptive]. More recently, transformer-based architectures have been used in reconstructing density matrices from measurement data [zhong2022quantum, palmieri2024enhancing, wei2024neural, ma2025tomography]. While these methods achieve remarkable accuracy, they often focus on recovering the density matrix rather than directly targeting scalar quantities. This incurs unnecessary computational overhead when only a few properties are of interest.
This work develops a Bayesian inference framework enhanced by residual learning and implemented with a PI neural network architecture, aiming to estimate scalar functions of an unknown quantum state from classical measurement data. Rather than reconstructing the full density matrix, our method focuses on efficient, scalable estimation of quantities such as fidelity and Rényi entanglement entropy. It exploits the unordered nature of non-adaptive measurements to design an architecture that generalizes for various system sizes as well as the numbers of measurements.
@@ III Problem setup
In a fully Bayesian approach, one would instead seek the posterior expectation of $f(\rho)$ given the observed data. That is, the ideal Bayesian estimator is the conditional expectation of $f(\rho)$ under the posterior distribution of $\rho$ after observing all $N$ measurement outcomes. Computing this exact posterior expectation is generally intractable for large quantum systems due to two fundamental challenges. First, it requires performing continuous integrations over state spaces. Second, the dimension of the quantum state $\rho$ grows exponentially with the number of qubits. In the next section, we describe our learning-based framework for approximating the Bayesian estimator, which enables the incorporation of arbitrary priors without explicitly evaluating a high-dimensional integral.
@@ IV.1 PI network and residual learning
the Bayesian posterior expectation of $f(\rho)$ using the measurement outcomes as input. Instead of directly predicting $f(\rho)$, we find it advantageous to leverage any existing estimator $g$ (such as the classical shadows estimator) and train the network to predict the residual error relative to this estimator. In other words, the network learns to predict the difference between the current estimator’s output $\hat{f}(\rho)$ and the true value $f(\rho)$. By correcting this difference, the network effectively learns to output a refined estimate closer to the ideal Bayesian value. That is, we train the model to predict the expected estimation error conditioned on all $N$ measurement results.
Even if each measurement choice can depend on all previous outcomes, the final posterior depends only on the multiset of observed values. Formally, we have the following observation.
@@
Our model is specifically designed to capture this permutation invariance. We employ the set transformer architecture [lee2019set], a neural network tailored for set-input problems, to process the collection of measurement outcome features. The set transformer uses a self-attention mechanism to produce a PI representation of the input set. They have been shown to be universal approximators for PI functions on sets. Following its original implementation, the network processes the input
@@
$\sigma(x,F)=F_{\text{c}}\times(1-\tanh|x|)+\mathbf{1}_{\mathbb{R}^{+}}(x)\tanh x,$
where $F_{\text{c}}=\max(\min(F,f_{\text{u}}),f_{\text{l}})$ is a clipped estimate and $\mathbf{1}_{A}(x)=[x\in A]$ is the indicator function. At $x=0$, we have $\sigma(0,F)=F_{\text{c}}$.
Alternatively, one could learn the estimator directly in the form $\text{Sigmoid}(x)$, rather than using a residual correction. However, as shown in the experimental results, this direct approach is outperformed by classical shadows as the number of measurements $N$ increases. This is because inverting the shadow requires learning parity-like functions that conventional learning algorithms notoriously struggle to approximate [bengio2005curse, eldan2016power, telgarsky2016benefits, daniely2020learning].
@@ IV.3 Encoding of measurement statistics
Each Clifford circuit is specified by its stabilizer tableau of size $2n\times(2n+1)$ [dehaene2003clifford, aaronson2004improved, bravyi2019simulation]. Concretely, each of the $2n$ rows consists of:
@@ IV.4 Application to fidelity estimation
As a first demonstration, we consider direct fidelity estimation (DFE), the task of estimating the fidelity of the unknown state $\rho$ with a target entangled state. Specifically, we take the target state to be the $n$-qubit Greenberger-Horne-Zeilinger (GHZ) state [greenberger1989going, greenberger1990bell] defined as
@@
The fidelity with the GHZ state is $f(\rho)=\langle\psi_{\mathrm{GHZ}}|\rho|\psi_{\mathrm{GHZ}}\rangle$. There exists an efficient procedure for estimating the fidelity of an unknown quantum state with the GHZ state [cha2025efficient]. Algorithm 2 in Appendix G describes this procedure. This baseline procedure is referred to as the “Shadow” estimator for fidelity. The estimator in Algorithm 2 has the following desirable properties.
@@ IV.5 Application to entanglement entropy estimation
As a second example, we consider the task of estimating a non-linear function of the state: the second-order Rényi entanglement entropy across a bipartition of the qubits [hastings2010measuring, islam2015measuring, brydges2019probing]. Let the qubits be partitioned into subset $A$ and its complement $B$ (for instance, $A$ could be the first $n/2$ qubits and $B$ the remaining $n/2$ qubits). The quantity of interest can be expressed as
@@ V.3 Second-order Rényi entanglement entropy estimation for the GHZ state
For Rényi entropy, each instance consists of a random $n$-qubit state (with a chosen bipartition $A|B$) along with $N$ simulated measurement outcomes according to Algorithm 3, and the corresponding ground truth entanglement entropy $f(\rho)=\text{tr}((\rho\otimes\rho)S_{A})$. The input state was the GHZ state, but an adversary randomly replaces it with a spurious state $\omega$, sampled according to the Hilbert-Schmidt metric [zyczkowski2001induced]. The replacement occurred with probability $p$ drawn uniformly from the interval $(0,1)$:
@@ V.4 Performance evaluation
FIG. 6(a) shows that direct learning of the target functional outperforms residual learning for small $N$. In this regime, the variance of the classical shadow estimator is so large that a crude estimate drawn from the feasible range of $f(\rho)$ may produce a better result. However, as $N$ increases and the shadow estimator becomes sufficiently accurate, residual learning outperforms direct learning. This indicates that approximating the shadow inversion from measurement data to a point estimate of $f(\rho)$ is much more challenging than learning only the bias of an existing estimator, since many shadow inversion steps include evaluations of parity-like functions (see Algorithm 2 and Eq. (11)).
@@ VI Discussion
Our approach inherits a key limitation of data-driven methods; since the network is trained to learn the posterior correction of the classical shadow estimator, it requires a large training set in order to generalize reliably. Each training sample must be labeled with an almost exact value of the target property—i.e. the ground-truth posterior mean—which in practice can only be obtained by performing a very large number of measurements. Thus, although our architecture scales efficiently with the number of qubits and measurements, the cost of generating labels poses an obstacle when the variance of the shadow estimator grows with system size.
The problem of estimating properties of a quantum state in the presence of imperfect control and readout [chen2021robust, koh2022classical, wu2024error] can be understood naturally within a Bayesian inference framework; one treats the parameters characterizing noisy unitaries and measurement channels as learnable latent variables, and then incorporates that information into posterior updates over the true quantum state.
Looking forward, one promising direction is to bypass the intermediate classical shadow reconstruction and instead update a posterior distribution directly over the target functional $f(\rho)$. By framing the problem as inference on the scalar quantity $f(\rho)$, one could potentially further reduce sample complexity and computational overhead. Another avenue for improvement lies in adaptive measurement strategies, where each subsequent measurement choice depends on previous outcomes to maximize information gain about $\rho$ or $f(\rho)$ [quek2021adaptive, lange2023adaptive]. Adaptivity can offer performance gains, and permutationally invariant architectures can still be employed to process measurement records.
@@ Appendix B Classical shadows
In the classical shadows protocol [huang2020predicting], one performs randomized measurements on many independent copies of an unknown state $\rho$. Each measurement outcome is used to construct a snapshot $\hat{\rho}$—an unbiased estimator of the state—such that by averaging these snapshots one can estimate any linear function of $\rho$. For example, for a Hermitian observable $O$, the classical shadows estimator satisfies $\mathbb{E}[\text{tr}(\hat{\rho}O)]=\text{tr}(\rho O)$, assuming $\mathbb{E}[\hat{\rho}]=\rho$ by construction.
@@ 2 $\texttt{sum}\leftarrow 0$
Algorithm 1 Classical shadow estimation of $\text{tr}(\rho O)$ [huang2020predicting]
@@
where each $U_{i}$ is independently sampled uniformly at random from the set $\{\mathbb{I},H,HS^{\dagger}\}$. This corresponds to performing local measurements in a randomly chosen basis from $\{X,Y,Z\}$. As will be discussed in the main text, this work considers DFE with Pauli measurements, where the target state is the GHZ state [greenberger1989going, greenberger1990bell]. In such cases, more efficient alternatives to the direct use of (5) are available, and we adopt one of these optimized methods.
@@
where $U$ is sampled uniformly from the $n$-qubit Clifford group $\mathrm{Cl}(2^{n})$ [nielsen2010quantum].